# Telegrapher’s Generative Model via Kac Flows

### v6

@@ 3 Motivation
and with the velocity field induced by $\alpha$, see [47, Example 4.12],
@@
where $Z_{t}$ is the normalization constant. There are a variety of approaches for approximating the vector field corresponding to the linear interpolation, see, e.g., [32, nüsken2024steintransportbayesianinference] for kernel-based methods. However, for certain asymmetric constellations of prior and target measure, a mode collapse of the neural model was numerically observed in [máté2023learning]. Indeed, it was analytically shown in [8] that for the Laplacian $p_{0}\propto e^{-|x|}$ centered at 0 and a Laplacian-like target distribution $p_{1}\propto e^{-2\min\{|x|,|x-m|\}}$ with a mode at 0 and a second one at $m>0$, the velocity norm grows exponentially with $m$, i.e. $\|v_{1}\|_{L_{2}(p_{1})}^{2}\gtrsim e^{\alpha m}$. Using Gaussian mixtures instead of Laplacians, an explosion of the type $\int_{0}^{1}\|v_{t}\|_{L_{2}(p_{t})}^{2}\,\mathrm{d}t\gtrsim m^{4}e^{\alpha m^{2}}$ was proven in [15, Proposition 9].
@@
Similarly to the diffusion case, the explosion (3.14) at heights $z$ close to the data ($z=0$) makes suitable truncations necessary when approximating the velocity field by a neural network, see [48, Ch. 3.2].
@@ 4.2 Relation of Kac’ Flow to 1D Telegrapher’s Equation
By the next lemma, there is an analytic representation of the distributional solution of the telegrapher’s equation (4). We refer to [31] and to [34] for the proof via distributions. For a characterization of measures as (tempered) distributions, see also [38, Chapter 4.4].
@@
The measure $\mu_{t}$ is exactly the probability distribution of the Kac process $K(t)$ starting in 0 from (4.6), see [20, Section 2].
@@
By [20, Section 2], the absolute continuous part of the distribution of the initially right-moving and initially left-moving Kac process $\pm S(t)$ starting in 0 is given by
@@ 5.1 Kac’s Insertion Method
This claim was again conjectured in [22] and a proof was given in [14, Theorem 4] using semigroup theory. $\diamond$
@@
In 1D, the solution of (5.1) is given by $w_{c}(t,x)=\frac{1}{2}(f_{0}(x+ct)+f_{0}(x-ct))$, so that (5.3) reproduces (4.8). Unfortunately, in higher dimensions, the telegrapher’s PDE solution (5.3) might not stay a probability density anymore. For example, the mass $\int_{\mathbb{R}^{3}}u(t,x)\,\mathrm{d}x$ in 3D behaves like $1-e^{-t}$, see [44, Ch. 2.2]. $\diamond$
@@
The statement can be found in [20], and a proof was given in [14, Theorem 5]. $\diamond$
@@
It is well-known that the Laplacian $\Delta:H^{2}(\mathbb{R}^{d})\subset L_{2}(\mathbb{R}^{d})\to L_{2}(\mathbb{R}^{d})$ is a self-adjoint linear operator with non-positive spectrum $\sigma(\Delta)=(-\infty,0]$. By [29, Corollary 2] the solution of (5.1) with $c=1$ is given by $w(t,\cdot)=C(t)f_{0}$, where $(C(t))_{t\geq 0}$ is a uniformly bounded so-called cosine operator function with $\|C(t)\|_{L_{2}\to L_{2}}\leq 1$ for all $t\geq 0$. In particular, it follows that $\|w(t,\cdot)\|_{L_{2}(\mathbb{R}^{d})}\leq\|f_{0}\|_{L_{2}(\mathbb{R}^{d})}$ is uniformly bounded in $t$. Now, [20, Theorem 2] yields the claim. ∎
@@ 6 Velocity Field of the Multi-Dimensional Kac Flow
The following general lemma states that the velocity field of a multi-dimensional Wasserstein flow can be retrieved by its 1D components if the probability flow admits a multiplicative decomposition (by independence). Related factorization results appear in the generator matching framework [18, Proposition 4] and in the discrete flow matching framework [25, Section 7.5]. These approaches consider probability evolutions arising from Markovian dynamics or from discrete flow formulations. Since the Kac process (4.6) is a non-Markovian [^5] Yet, it becomes Markovian by considering the joint process $(K(t),\dot{K}(t))$. process [12, Chapter 12] in a continuous setting, and for completeness, we provide a rigorous and general proof here.
@@
More precisely, we can test above equations by product test functions $\phi(t,x)=\prod_{i=1}^{d}\phi^{i}(t,x^{i})$ with $\phi^{i}\in C_{c}^{\infty}((0,\infty)\times\mathbb{R})$, and obtain the result in the weak formulation. Finally, since the linear subspace generated by such products is dense in $C_{c}^{\infty}((0,\infty)\times\mathbb{R}^{d})$ and by Lebesgue’s dominated convergence, this yields the claim. We refer to Section A.2 in the appendix for the full mathematical details. ∎
@@
By Proposition 5.6, $\mu_{t}(\cdot|x_{0})$ admits a velocity field $v_{t}(\cdot|x_{0})$ satisfying the continuity equation (2.2). By construction, the components $(X_{t}^{1},...,X_{t}^{d})$ of $\mathbf{X}_{t}$ are mutually independent. Hence, $\mu_{t}(x|x_{0})=\prod_{i=1}^{d}\mu_{t}^{i}(x^{i}|x_{0}^{i})$, where the one-dimensional probability flows $\mu_{t}^{i}(\cdot|x_{0}^{i})$ fulfill a continuity equation with 1D velocity fields $v_{t}^{i}(\cdot|x_{0}^{i})$. By Lemma 6.1, it holds the decomposition (6.4), and Theorem 4.4 immediately gives the formula (6.9). Finally, [27, Theorem 3.3], resp. [47, Theorem 6.3] provides (6.10). ∎
@@
Having identified the conditional vector field $v_{t}(\cdot|x_{0})$ for the conditional distribution $\mu_{t}(\cdot|x_{0})$ of the Kac process, we get the following result by employing [24], see also [47, Theorem 5.1, Remark 5.3], with slight adjustments to our setting.
@@
where $\mathcal{U}(0,T)\times_{t}u_{t}$ is understood in the sense of disintegration, see [47, Ch. 2.3], but here we have no direct access to $v$. However, following the same lines as in [24], we see that
@@ 7.2 Image Generation
In Table 1 we report the FID results of our Kac methods via $M_{t}$ as well as the diffusion baselines via $R_{t}$. We implemented the baseline models ourselves and denote them by "Diff" for $g(t)=t$ and "FM" for $g(t)=t^{2}$. In Figures 6, 7 we plot a selection of the generated images. Recall that the parameter $a$ corresponds to the damping coefficient and $c$ to the propagation speed. The schedule $g(t)=t$ generally led to worse performance, especially for the baseline and our methods with diffusion-like parameters. However certain combinations of high damping and low velocity pairs $(a,c)$ such as $(25,2)$ were able to perform well. In contrast the schedule $g(t)=t^{2}$ improved the baseline substantially, and our methods across the board. With this experiment we aim to demonstrate the scalability of our method, an extensive study of the effect of the parameters $(a,c)$ is beyond the scope of the paper and left to future work. For further visual examples and a nearest neighbour analysis, see Appendix A.5.
@@
Table 1: FID (100-step Euler) for $(a,c)$-Kac under two time schedules.
@@ Acknowledgement.
J. Chemseddine acknowledges funding of the DFG project STE 571/17-2 within the SPP 2298 "Theoretical Foundations of Deep Learning". The work of R. Duong, P. Friz and G. Steidl is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – The Berlin Mathematics Research Center MATH+ (EXC-2046/1, EXC-2046/2, project ID: 390685689). The work of P. Friz is also supported by DFG CRC/TRR 388, project A02. Our thanks go to Gregor Kornhardt, Matthias Liero, Nicolaj Rux, and Weiqiao Han for fruitful discussions. We are especially grateful to Weiqiao Han for correcting an error in our implementation.
@@ A.1 Lipschitz property of the Mean-reverting Kac process
Define the tensor product space $C_{\mathrm{c},+}^{\infty}((0,T))\otimes\bigotimes_{i=1}^{d}C_{\mathrm{c}}^{\infty}(\mathbb{R})$ as the linear span of functions
@@
with functions $\psi_{0}\in C_{\mathrm{c}}^{\infty}((0,T)),\,\psi_{i}\in C_{\mathrm{c}}^{\infty}(\mathbb{R}),\,i=1,\dots,d$, such that $\psi_{0}$ is nonnegative, and $\psi_{0}^{\frac{1}{d}}\in C_{\mathrm{c}}^{\infty}((0,T))$.
By the following result, this subspace is dense in $C_{\mathrm{c}}^{\infty}((0,T)\times\mathbb{R}^{d})$.
@@
The space $C_{\mathrm{c},+}^{\infty}((0,T))\otimes\bigotimes_{i=1}^{d}C_{\mathrm{c}}^{\infty}(\mathbb{R})$ is dense in $C_{\mathrm{c}}^{\infty}((0,T)\times\mathbb{R}^{d})$ in the following sense: for all $\psi\in C_{\mathrm{c}}^{\infty}((0,T)\times\mathbb{R}^{d})$, there exists a compact set $K\subset(0,T)\times\mathbb{R}^{d}$ containing $\mathrm{supp}\,\psi$, and a sequence $(\psi_{n})\subset C_{\mathrm{c},+}^{\infty}((0,T))\otimes\bigotimes_{i=1}^{d}C_{\mathrm{c}}^{\infty}(\mathbb{R})$ compactly supported on $K$, such that $D^{\alpha}\psi_{n}$ converges uniformly to $D^{\alpha}\psi$ on $K$ for any $\alpha\in\mathbb{N}_{0}^{d+1}$.
@@
Now, consider nonnegative bump functions $\xi_{0},\dots,\xi_{d}\in C_{\mathrm{c}}^{\infty}(\mathbb{R})$ such that
@@
Defining $K_{i}\coloneqq\mathrm{supp}\,\xi_{i},~i=0,\dots,d$, we see that $K\coloneqq K_{0}\times\dots\times K_{d}$ is a compact set containing $\mathrm{supp}\,\psi$. Note that $\xi_{0}$ can be chosen such that $K_{0}\subset(0,T)$, i.e. $\xi_{0}\in C_{\mathrm{c}}^{\infty}((0,T))$, and such that $\xi_{0}$ has smooth $d$-th roots $\xi_{0}^{\frac{1}{d}}\in C_{\mathrm{c}}^{\infty}((0,T))$. [^8] Pick a nonnegative bump function $\tilde{\xi}_{0}\in C_{\mathrm{c}}^{\infty}((0,T))$ fulfilling (A.12), and consider $\xi_{0}\coloneqq(\tilde{\xi}_{0})^{d}$.
@@
$\psi_{n}(t,x)\coloneqq\xi_{0}(t)\xi_{1}(x_{1})\dots\xi_{d}(x_{d})p_{n}(t,x),\quad t\in(0,T),\,x=(x_{1},\dots,x_{d})\in\mathbb{R}^{d}.$ (A.13)
@@
with $m_{n}$ being the degree of $p_{n}$. By the positivity of $t\in(0,T)$ and the choice of $\xi_{0}$, the function $\psi_{0}(t)\coloneqq t^{j}\xi_{0}(t)$ has smooth $d$-th root $\psi_{0}^{\frac{1}{d}}\in C_{\mathrm{c}}^{\infty}((0,T))$ for any $j\in\mathbb{N}_{0}$. Hence, the functions $\psi_{n}$ are linear combinations of tensor product functions (A.10), i.e. $\psi_{n}\in C_{\mathrm{c},+}^{\infty}((0,T))\otimes\bigotimes_{i=1}^{d}C_{\mathrm{c}}^{\infty}(\mathbb{R})$. By construction, $\psi_{n}$ is compactly supported on $K$.
@@
Remark A.3.
Note that the requirement $\psi_{0}^{\frac{1}{d}}\in C_{\mathrm{c}}^{\infty}((0,T))$ is not automatically given by the smoothness of $\psi_{0}\in C_{\mathrm{c}}^{\infty}((0,T))$. For example, take $\psi_{0}(t)\coloneqq t^{2}$ on $(-\frac{1}{2},\frac{1}{2})$ and smoothly cut it off to obtain a nonnegative function $\psi_{0}\in C_{\mathrm{c}}^{\infty}((-1,1))$. Then, $\psi_{0}(t)^{\frac{1}{2}}=|t|$ on $(-\frac{1}{2},\frac{1}{2})$ is not smooth. Even assuming the strict positivity of $\psi_{0}$ on its support is not enough: in [13], see also [5, Section 2], an example is given via
$\psi_{0}(t)\coloneqq\begin{cases}e^{-\frac{1}{t}}(\sin^{2}(\frac{\pi}{t})+e^{-\frac{1}{t}}),~&t>0,\\ 0,~&t\leq 0.\end{cases}$ (A.15)
By smoothly cutting it off at e.g. $t=1/2$, we can obtain a smooth function $\psi_{0}\in C_{\mathrm{c}}^{\infty}((-1,1))$, which is strictly positive on its support. But the square root $\psi_{0}^{\frac{1}{2}}$ is not $C^{2}$ due to strong oscillations at $t=0$. Naturally, we obtain counter-examples in $C_{\mathrm{c}}^{\infty}((0,T))$ by translation. $\diamond$
Now, consider tensor products of the following form
$\psi:(0,T)\times\mathbb{R}^{d}\to\mathbb{R},\quad(t,x_{1},\dots,x_{d})\mapsto\prod_{i=1}^{d}\psi_{i}(t,x_{i}),$ (A.16)
with functions $\psi_{i}\in C_{\mathrm{c}}^{\infty}((0,T)\times\mathbb{R}),\,i=1,\dots,d$.
Corollary A.4.
The linear span of functions (A.16) is dense in $C_{\mathrm{c}}^{\infty}((0,T)\times\mathbb{R}^{d})$ in the sense given in Lemma (A.2).
Proof.
Just note that any functions $\psi$ given by (A.10) can be written as
$\psi(t,x)=(\psi_{0}(t)^{\frac{1}{d}})^{d}\prod_{i=1}^{d}\psi_{i}(x_{i})=\prod_{i=1}^{d}\tilde{\psi}_{i}(t,x_{i}),$ (A.17)
where $\tilde{\psi}_{i}(t,x_{i})\coloneqq\psi_{0}(t)^{\frac{1}{d}}\psi_{i}(x_{i})$. By construction, we have $\psi_{0}^{\frac{1}{d}}\in C_{\mathrm{c}}^{\infty}((0,T))$, and thus $\tilde{\psi}_{i}\in C_{\mathrm{c}}^{\infty}((0,T)\times\mathbb{R})$, i.e. $\psi$ is of the form (A.16). Now, the claim follows by Lemma A.2. ∎
@@
Take any smooth test function $\psi$ of the form (A.16). Then, we have by the product rule that
$\int_{0}^{T}\int_{\mathbb{R}^{d}}\partial_{t}\psi(t,x)\,\mathrm{d}\mu_{t}(x)\,\mathrm{d}{t}=\int_{0}^{T}\int_{\mathbb{R}^{d}}\sum_{i=1}^{d}\partial_{t}\psi_{i}(t,x_{i})\prod_{j\neq i}^{d}\psi_{j}(t,x_{j})\,\mathrm{d}\mu_{t}(x)\,\mathrm{d}{t}.$ (A.18)
Using the assumptions (6.1), (6.2) and Fubini’s theorem yields
$=\sum_{i=1}^{d}\int_{\mathbb{R}^{d-1}}\left(\int_{0}^{T}\int_{\mathbb{R}}\partial_{t}\psi_{i}(t,x_{i})\,\mathrm{d}\mu^{i}_{t}(x_{i})\,\mathrm{d}{t}\right)\prod_{j\neq i}^{d}\psi_{j}(t,x_{j})\,\mathrm{d}\mu_{t}^{j}(x_{j})$ (A.19)
$=\sum_{i=1}^{d}\int_{\mathbb{R}^{d-1}}\left(-\int_{0}^{T}\int_{\mathbb{R}}v_{t}^{i}(x_{i})\frac{\,\mathrm{d}}{\,\mathrm{d}x_{i}}\psi_{i}(t,x_{i})\,\mathrm{d}\mu^{i}_{t}(x_{i})\,\mathrm{d}{t}\right)\prod_{j\neq i}^{d}\psi_{j}(t,x_{j})\,\mathrm{d}\mu_{t}^{j}(x_{j})$ (A.20)
$=\sum_{i=1}^{d}\int_{\mathbb{R}^{d-1}}\left(-\int_{0}^{T}\int_{\mathbb{R}}v_{t}^{i}(x_{i})\frac{\partial}{\partial x_{i}}\psi(t,x)\,\mathrm{d}\mu^{i}_{t}(x_{i})\,\mathrm{d}{t}\right)\prod_{j\neq i}^{d}\,\mathrm{d}\mu_{t}^{j}(x_{j})$ (A.21)
@@
where the vector field $v_{t}$ is given by (6.3). Note that $v_{t}$ satisfies the integrability (2.3) by the assumptions $\mu^{i}_{t}\in AC^{2}((0,T);\allowbreak\mathcal{P}_{2}(\mathbb{R}))$. Hence, above equations also hold true for any linear combination of test functions $\psi$ given by (A.16). Finally, by Corollary A.4 and Lebesgue’s dominated convergence we infer the continuity equation in the variational formulation (A.9) for all $\psi\in C_{\mathrm{c}}^{\infty}((0,T)\times\mathbb{R}^{d})$. ∎

### v7

@@ 3 Motivation
and with the velocity field induced by $\alpha$, see [46, Example 4.12],
@@
where $Z_{t}$ is the normalization constant. There are a variety of approaches for approximating the vector field corresponding to the linear interpolation, see, e.g., [31, nüsken2024steintransportbayesianinference] for kernel-based methods. However, for certain asymmetric constellations of prior and target measure, a mode collapse of the neural model was numerically observed in [máté2023learning]. Indeed, it was analytically shown in [8] that for the Laplacian $p_{0}\propto e^{-|x|}$ centered at 0 and a Laplacian-like target distribution $p_{1}\propto e^{-2\min\{|x|,|x-m|\}}$ with a mode at 0 and a second one at $m>0$, the velocity norm grows exponentially with $m$, i.e. $\|v_{1}\|_{L_{2}(p_{1})}^{2}\gtrsim e^{\alpha m}$. Using Gaussian mixtures instead of Laplacians, an explosion of the type $\int_{0}^{1}\|v_{t}\|_{L_{2}(p_{t})}^{2}\,\mathrm{d}t\gtrsim m^{4}e^{\alpha m^{2}}$ was proven in [14, Proposition 9].
@@
Similarly to the diffusion case, the explosion (3.14) at heights $z$ close to the data ($z=0$) makes suitable truncations necessary when approximating the velocity field by a neural network, see [47, Ch. 3.2].
@@ 4.2 Relation of Kac’ Flow to 1D Telegrapher’s Equation
By the next lemma, there is an analytic representation of the distributional solution of the telegrapher’s equation (4). We refer to [30] and to [33] for the proof via distributions. For a characterization of measures as (tempered) distributions, see also [37, Chapter 4.4].
@@
The measure $\mu_{t}$ is exactly the probability distribution of the Kac process $K(t)$ starting in 0 from (4.6), see [19, Section 2].
@@
By [19, Section 2], the absolute continuous part of the distribution of the initially right-moving and initially left-moving Kac process $\pm S(t)$ starting in 0 is given by
@@ 5.1 Kac’s Insertion Method
This claim was again conjectured in [21] and a proof was given in [13, Theorem 4] using semigroup theory. $\diamond$
@@
In 1D, the solution of (5.1) is given by $w_{c}(t,x)=\frac{1}{2}(f_{0}(x+ct)+f_{0}(x-ct))$, so that (5.3) reproduces (4.8). Unfortunately, in higher dimensions, the telegrapher’s PDE solution (5.3) might not stay a probability density anymore. For example, the mass $\int_{\mathbb{R}^{3}}u(t,x)\,\mathrm{d}x$ in 3D behaves like $1-e^{-t}$, see [43, Ch. 2.2]. $\diamond$
@@
The statement can be found in [19], and a proof was given in [13, Theorem 5]. $\diamond$
@@
It is well-known that the Laplacian $\Delta:H^{2}(\mathbb{R}^{d})\subset L_{2}(\mathbb{R}^{d})\to L_{2}(\mathbb{R}^{d})$ is a self-adjoint linear operator with non-positive spectrum $\sigma(\Delta)=(-\infty,0]$. By [28, Corollary 2] the solution of (5.1) with $c=1$ is given by $w(t,\cdot)=C(t)f_{0}$, where $(C(t))_{t\geq 0}$ is a uniformly bounded so-called cosine operator function with $\|C(t)\|_{L_{2}\to L_{2}}\leq 1$ for all $t\geq 0$. In particular, it follows that $\|w(t,\cdot)\|_{L_{2}(\mathbb{R}^{d})}\leq\|f_{0}\|_{L_{2}(\mathbb{R}^{d})}$ is uniformly bounded in $t$. Now, [19, Theorem 2] yields the claim. ∎
@@ 6 Velocity Field of the Multi-Dimensional Kac Flow
The following general lemma states that the velocity field of a multi-dimensional Wasserstein flow can be retrieved by its 1D components if the probability flow admits a multiplicative decomposition (by independence). Related factorization results appear in the generator matching framework [17, Proposition 4] and in the discrete flow matching framework [24, Section 7.5]. These approaches consider probability evolutions arising from Markovian dynamics or from discrete flow formulations. Since the Kac process (4.6) is a non-Markovian [^5] Yet, it becomes Markovian by considering the joint process $(K(t),\dot{K}(t))$. process [12, Chapter 12] in a continuous setting, and for completeness, we provide a rigorous and general proof here.
@@
More precisely, we can test above equations with product test functions $\phi(t,x)=\phi_{0}(t)\prod_{i=1}^{d}\phi_{i}(x^{i})$, where $\phi_{i}\in C_{c}^{\infty}(\mathbb{R})$, and obtain the result in the weak formulation. Finally, since the linear subspace generated by such products is dense in $C_{c}^{\infty}((0,\infty)\times\mathbb{R}^{d})$ and by Lebesgue’s dominated convergence, this yields the claim. We refer to Section A.2 in the appendix for the full mathematical details. ∎
@@
By Proposition 5.6, $\mu_{t}(\cdot|x_{0})$ admits a velocity field $v_{t}(\cdot|x_{0})$ satisfying the continuity equation (2.2). By construction, the components $(X_{t}^{1},...,X_{t}^{d})$ of $\mathbf{X}_{t}$ are mutually independent. Hence, $\mu_{t}(x|x_{0})=\prod_{i=1}^{d}\mu_{t}^{i}(x^{i}|x_{0}^{i})$, where the one-dimensional probability flows $\mu_{t}^{i}(\cdot|x_{0}^{i})$ fulfill a continuity equation with 1D velocity fields $v_{t}^{i}(\cdot|x_{0}^{i})$. By Lemma 6.1, it holds the decomposition (6.4), and Theorem 4.4 immediately gives the formula (6.9). Finally, [26, Theorem 3.3], resp. [46, Theorem 6.3] provides (6.10). ∎
@@
Having identified the conditional vector field $v_{t}(\cdot|x_{0})$ for the conditional distribution $\mu_{t}(\cdot|x_{0})$ of the Kac process, we get the following result by employing [23], see also [46, Theorem 5.1, Remark 5.3], with slight adjustments to our setting.
@@
where $\mathcal{U}(0,T)\times_{t}u_{t}$ is understood in the sense of disintegration, see [46, Ch. 2.3], but here we have no direct access to $v$. However, following the same lines as in [23], we see that
@@ 7.2 Image Generation
In Table 7.1 we report the FID results of our Kac methods via $M_{t}$ as well as the diffusion baselines via $R_{t}$. We implemented the baseline models ourselves and denote them by "Diff" for $g(t)=t$ and "FM" for $g(t)=t^{2}$. In Figures 6, 7 we plot a selection of the generated images. Recall that the parameter $a$ corresponds to the damping coefficient and $c$ to the propagation speed. The schedule $g(t)=t$ generally led to worse performance, especially for the baseline and our methods with diffusion-like parameters. However certain combinations of high damping and low velocity pairs $(a,c)$ such as $(25,2)$ were able to perform well. In contrast the schedule $g(t)=t^{2}$ improved the baseline substantially, and our methods across the board. With this experiment we aim to demonstrate the scalability of our method, an extensive study of the effect of the parameters $(a,c)$ is beyond the scope of the paper and left to future work. For further visual examples and a nearest neighbour analysis, see Appendix A.5.
@@
Table 7.1: FID (100-step Euler) for $(a,c)$-Kac under two time schedules.
@@ Acknowledgement.
J. Chemseddine acknowledges funding of the DFG project STE 571/17-2 within the SPP 2298 "Theoretical Foundations of Deep Learning". The work of R. Duong, P. Friz and G. Steidl is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – The Berlin Mathematics Research Center MATH+ (EXC-2046/1, EXC-2046/2, project ID: 390685689). The work of P. Friz is also supported by DFG CRC/TRR 388, project A02. Our thanks go to Gregor Kornhardt, Matthias Liero, and Nicolaj Rux for fruitful discussions. We are especially grateful to Weiqiao Han for correcting an error in our implementation, and to Flinn Raffael for correcting a mistake in the proof of Lemma 6.1.
@@ A.1 Lipschitz property of the Mean-reverting Kac process
Define the tensor product space $C_{\mathrm{c}}^{\infty}((0,T))\otimes\bigotimes_{i=1}^{d}C_{\mathrm{c}}^{\infty}(\mathbb{R})$ as the linear span of functions
@@
with functions $\psi_{0}\in C_{\mathrm{c}}^{\infty}((0,T)),\,\psi_{i}\in C_{\mathrm{c}}^{\infty}(\mathbb{R}),\,i=1,\dots,d$.
By the following well-known result, this subspace is dense in $C_{\mathrm{c}}^{\infty}((0,T)\times\mathbb{R}^{d})$.
@@
The space $C_{\mathrm{c}}^{\infty}((0,T))\otimes\bigotimes_{i=1}^{d}C_{\mathrm{c}}^{\infty}(\mathbb{R})$ is dense in $C_{\mathrm{c}}^{\infty}((0,T)\times\mathbb{R}^{d})$ in the following sense: for all $\psi\in C_{\mathrm{c}}^{\infty}((0,T)\times\mathbb{R}^{d})$, there exists a compact set $K\subset(0,T)\times\mathbb{R}^{d}$ containing $\mathrm{supp}\,\psi$, and a sequence $(\psi_{n})\subset C_{\mathrm{c}}^{\infty}((0,T))\otimes\bigotimes_{i=1}^{d}C_{\mathrm{c}}^{\infty}(\mathbb{R})$ compactly supported on $K$, such that $D^{\alpha}\psi_{n}$ converges uniformly to $D^{\alpha}\psi$ on $K$ for any $\alpha\in\mathbb{N}_{0}^{d+1}$.
@@
Now, consider bump functions $\xi_{0},\dots,\xi_{d}\in C_{\mathrm{c}}^{\infty}(\mathbb{R})$ such that
@@
Defining $K_{i}\coloneqq\mathrm{supp}\,\xi_{i},~i=0,\dots,d$, we see that $K\coloneqq K_{0}\times\dots\times K_{d}$ is a compact set containing $\mathrm{supp}\,\psi$. Note that $\xi_{0}$ can be chosen such that $K_{0}\subset(0,T)$, i.e. $\xi_{0}\in C_{\mathrm{c}}^{\infty}((0,T))$.
@@
$\psi_{n}(t,x)\coloneqq\xi_{0}(t)\xi_{1}(x_{1})\dots\xi_{d}(x_{d})p_{n}(t,x),\quad t\in(0,T),~x=(x_{1},\dots,x_{d})\in\mathbb{R}^{d}.$ (A.13)
@@
with $m_{n}$ being the degree of $p_{n}$. Hence, the functions $\psi_{n}$ are linear combinations of tensor product functions (A.10), i.e. $\psi_{n}\in C_{\mathrm{c}}^{\infty}((0,T))\otimes\bigotimes_{i=1}^{d}C_{\mathrm{c}}^{\infty}(\mathbb{R})$. By construction, $\psi_{n}$ is compactly supported on $K$.
@@
Furthermore, we need the following handy proposition, which can be found in [4, Proposition 16.3]. It allows for a more manageable treatment of the weak continuity equation (A.9).
Proposition A.3.
Let $\mu_{t}:(0,T)\to\mathcal{P}(\mathbb{R}^{d})$ be a narrowly continuous curve of probability measures and let $v_{t}$ be a vector field such that $v_{t}\in L_{1}(\mu_{t})$ and $\|v_{t}\|_{L_{1}(\mu_{t})}\in L_{1}(0,T)$. Then the following are equivalent:
  • (i)
    the curve $\mu_{t}$ solves (A.9).
  • (ii)
    for any $\psi\in C_{\mathrm{c}}^{\infty}(\mathbb{R}^{d})$, it holds that $t\mapsto\int_{\mathbb{R}^{d}}\psi(x)\,\mathrm{d}\mu_{t}(x)$ is absolutely continuous on $[0,T]$ and its derivative is $\int_{\mathbb{R}^{d}}\nabla\psi(x)\cdot v_{t}(x)\,\mathrm{d}\mu_{t}(x)$.
@@
Take any smooth test function $\psi$ of the form (A.10). Then, using the factorization assumption (6.1) we have
$\int_{0}^{T}\int_{\mathbb{R}^{d}}\partial_{t}\psi(t,x)\,\mathrm{d}\mu_{t}(x)\,\mathrm{d}{t}=\int_{0}^{T}\partial_{t}\psi_{0}(t)\prod_{i=1}^{d}\left(\int_{\mathbb{R}}\psi_{i}(x_{i})\,\mathrm{d}\mu_{t}^{i}(x_{i})\right)\,\mathrm{d}{t}.$ (A.15)
Since by assumption, the 1D flows $\mu_{t}^{i}$ satisfy the continuity equation (6.2) in the distributional sense, Proposition A.3, $(i)\Rightarrow(ii)$, yields that the functions $t\mapsto\int_{\mathbb{R}}\psi_{i}(x_{i})\,\mathrm{d}\mu_{t}^{i}(x_{i})$ are absolutely continuous with weak (and strong a.e.) derivative $\int_{\mathbb{R}}v_{t}^{i}(x_{i})\frac{\,\mathrm{d}}{\,\mathrm{d}x_{i}}\psi_{i}(x_{i})\,\mathrm{d}\mu_{t}^{i}(x_{i})$.
By integration by parts and the product rule for absolutely continuous functions, it follows
$=-\int_{0}^{T}\psi_{0}(t)~\partial_{t}\prod_{i=1}^{d}\left(\int_{\mathbb{R}}\psi_{i}(x_{i})\,\mathrm{d}\mu_{t}^{i}(x_{i})\right)\,\mathrm{d}{t}$ (A.16)
$=-\sum_{i=1}^{d}\int_{0}^{T}\psi_{0}(t)~\left(\int_{\mathbb{R}}v_{t}^{i}(x_{i})\frac{\,\mathrm{d}}{\,\mathrm{d}x_{i}}\psi_{i}(x_{i})\,\mathrm{d}\mu_{t}^{i}(x_{i})\right)\prod_{j\neq i}^{d}\left(\int_{\mathbb{R}}\psi_{j}(x_{j})\,\mathrm{d}\mu_{t}^{j}(x_{j})\right)\,\mathrm{d}{t}$ (A.17)
$=-\sum_{i=1}^{d}\int_{0}^{T}\int_{\mathbb{R}^{d}}v_{t}^{i}(x_{i})\frac{\partial}{\partial x_{i}}\psi(t,x)\,\mathrm{d}\mu_{t}(x)\,\mathrm{d}{t}$ (A.18)
@@
where the vector field $v_{t}$ is given by (6.3). Note that $v_{t}$ satisfies the integrability (2.3) by the assumptions $\mu^{i}_{t}\in AC^{2}((0,T);\allowbreak\mathcal{P}_{2}(\mathbb{R}))$. Hence, above equations also hold true for any linear combination of test functions $\psi$ given by (A.10). Finally, by Lemma A.2 and Lebesgue’s dominated convergence we infer the continuity equation in the variational formulation (A.9) for all $\psi\in C_{\mathrm{c}}^{\infty}((0,T)\times\mathbb{R}^{d})$. ∎